Please answer, i am confused (show work)
A line passes through points (-5,2) and (2,y) has a slope of 3. Find y.

Answer by Alan3354(69443)

Slope m = diffy/diffx = 3
(2-y)/(-5-2) = 3
(2-y)/-7 = 3
2-y = -21
y = 23